OpenStudy (anonymous):

2.Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (anonymous):

@radar help?

OpenStudy (anonymous):

@agent0smith help?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@wio

OpenStudy (agent0smith):

Sqrt(x)^4 is one.

OpenStudy (anonymous):

can you show me steps on how it simplified back down to x^2?

OpenStudy (agent0smith):

Sqrt(x)^4 = (x^1/2)^4 = x^4/2 = x^2

OpenStudy (anonymous):

can you do 2 more equations & show me?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

I'll do one, and you should be able to find one more using the same pattern First: Sqrt(x)^4 = (x^1/2)^4 = x^4/2 = x^2 Second: cuberoot(x)^6 = (x^1/3)^6 = x^6/3 = x^2

OpenStudy (anonymous):

My problem with this question: You say the number of steps matters, but haven't really defined what counts as a step.

OpenStudy (anonymous):

x*x*x/x

OpenStudy (anonymous):

one more? @agent0smith

OpenStudy (anonymous):

please ? @agent0smith

OpenStudy (agent0smith):

You have one more above, wio's

OpenStudy (anonymous):

I don't understand that one..

OpenStudy (anonymous):

could you give me one more?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

First: Sqrt(x)^4 = (x^1/2)^4 = x^4/2 = x^2 Second: cuberoot(x)^6 = (x^1/3)^6 = x^6/3 = x^2 Follow the pattern, find a third.

OpenStudy (anonymous):

can you please give me one more?/: I have like 40 assignments due tonight.

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

(x^1/2)^4 (x^1/3)^6 Look at the pattern, how the fractions and the numbers change. Copy and paste one of those and make a third.